OpenStudy (ttop0816):

7 · 15 can be written as _____. 5(7 + 10) 10(7 + 5) 7(10 + 5)

OpenStudy (anonymous):

7(10+5)

OpenStudy (anonymous):

please do not just give answers.

OpenStudy (calculusxy):

A easy trick can be to just find the product of 15*7. Then with that product see which product equals to the ones written on the answer choices.

OpenStudy (anonymous):

because 7*15=105 and 7(10+5)=105

OpenStudy (anonymous):

7 * 15 = 105 so: 5(7 + 10) 35 + 50 85 > NO 10(7 + 5) 70 + 50 120 > NO 7(10 + 5) 70 + 35 > 105 YES 105

OpenStudy (anonymous):

* typo * do not mind the 105 at the way bottom. lol sorry.